4. Fill in the blank spaces in the table. Units of good XTotal cost Marginal cost $40 $89 $149 2
Answer #1

ANSWER:

marginal cost = change in total cost / change in quantity

Units of good x Total cost Marginal cost
1 $40 = ($40 - $0) / (1 - 0) = $40
2 $89 = ($89 - $40) / (2 - 1) = $49
3 $149 = ($149 - $89) / (3 - 2) = $60